OpenStudy (anonymous):

How would you complete the square of x^2+10x+14=0

OpenStudy (anonymous):

bring 14 to the other side x^2+10x=-14 do (b/2)^2 which is 25 add 25 to both sides x^2+10x+25=11 factor (x+5)^2=11 square both side x+5= Sqrt 11 subtract 5 x= -5+-sqrt11

OpenStudy (anonymous):

thank you :)